In the reaction betweem H2O2 and I-, 15 mL of 0.30M KI, 69. mL of 0.10M H2O2 and 19 mL of 0.020M Na2S2O3 were mixed in a flask containing starch and buffer and the total volume was made up to 500mL with distilled water. The solution turned blue (indicating presence of I2) in 2.0 minutes. Calculate the intial rate of disappearance of H2O2 in units of M min-1 in 2 significant digits.
1
Expert's answer
2011-08-01T08:55:10-0400
Initial concentration of H2O2 in solution was:
C(H2O2) = C1 V1/V2 = 0.069 L × 0.1 M/0.5 L = 0.0138 M
So, initial rate is:
r= −d[H2O2]/dt = ΔC(H2O2)/Δt = 0.0138 M / 2 min = 0.0069 [M min−1]